# Instruction-Guided Scene Text Recognition

### v2

@@ 4.2 Ablations
Instruction variants. We conduct experiments to validate the effectiveness of the triplet instructions. The results in Tab. III include IGTR-PR (PR), IGTR-AR (AR) and IGTR-ER (ER). In (a), only recognition instructions are utilized for IGTR training. In subsequent variants denoted by (b), (c), (d), and (j), we add questions related to different attributes progressively. For these variants, the attribute prediction instructions are trained first, and then the model is fine-tuned on recognition instructions. The improvements reveal that each attribute contributes positively, supporting that understanding diverse character attributes aids in recognition. Comparing the use of full instructions in (j) to (a), PR improved by +1.94%, AR by +3.59%, and ER by +2.99%, indicating that AR and ER benefit more from the diverse questions, and meanwhile, making use of attributes in condition is helpful. On the other hand, we perform (e) with all questions but without condition ($cond$), and then progressively add a condition mode in (f)-(j), where the suffix is the condition index in Tab. I. IGTR enables AR and ER capabilities only when $cond$-3 and $cond$-4 are successively added. The results show that condition not only enhances accuracy but also realizes different recognition pipelines conveniently by adapting the instructions. Additionally, by increasing K, the number of partitions, the accuracy steadily improved, again providing intuitive confirmation of the importance of instruction richness and diversity.
@@
CMFF components. As shown in Fig. 2, CMFF incorporates four cross-attention stages for inter-modal interaction. We devise variants by removing different stages, and the results are presented in Tab. IV. Removing any stage all results in a decrease in accuracy. Notably, the most significant accuracy drop of 2.81% occurs when both query-to-condition and query-to-image stages are removed, emphasizing the necessity of sufficient feature interaction and fusion before making predictions. In addition, for AR and ER the condition is preceding decoded characters while for PR is set to None. It is anticipated that query-to-condition and query-to-image have a more pronounced impact on AR and ER, while PR is relatively less affected. The assumption explains a consistent accuracy degradation of 3.63% for AR and 2.95% for ER. Interestingly, PR also experiences a 1.66% decrease, albeit less severe than AR and ER. This could be attributed to that the query-to-condition and query-to-image stages have encoded ample instruction patterns during training, which are beneficial to all three recognition pipelines. In addition, PR, AR, and ER report inference speeds of 3.98ms, 10.3ms and 9.52ms. The speed differences are in accordance with their recognition pipelines. Meanwhile, they run faster than many previous models following similar recognition pipelines. Notably, as shown in Tab. IV, removing cross-attention stages only triggers minor speed acceleration. The results, in turn, verify that CMFF is quite efficient.
@@
TABLE VI: Zero-Shot recognition results of using different questions.
@@
IGTR-PD 63.88 84.93 88.79 72.25
IGTR-PD-TS 67.47 86.61 89.86 75.05
TABLE VIII: Morphologically similar characters and mis-recognitions on CTR [59]. Take “{CJK}UTF8gbsn莱-{CJK}UTF8gbsn菜” and 8/5 as the example. It means that without considering TS, “{CJK}UTF8gbsn莱” has been recognized as “{CJK}UTF8gbsn菜” for 8 times, and the number is reduced to 5 when TS is employed.
@@
TABLE X: Results on English benchmarks tested against existing models when trained on real-world Union14M-L dataset. denotes that the result is obtained by training the model on Union14M-L using the code they released.
@@
Scalability of IGTR. We conduct ablations on training data volume and model size. Specifically, we select IGTR-PR and vanilla PR, where the latter is trained solely using the PR instruction. The results are presented in the left half of Tab. V, where different training epochs are considered. We observe a notable accuracy gain of 2.78% when increasing epochs from 20 to 60 for IGTR-PR. This improvement is accompanied by the expansion of training data volume, as IGTR-PR samples attribute prediction instructions individually at each epoch. The trained model benefits from richer instructions. In contrast, the improvement of vanilla PR is 0.91%, solely from more training epochs that generate a slightly better model optimization. The significant improvement gap (2.78% v.s. 0.91%) implies that our instruction-guided training is more flexible and can produce more powerful STR models when necessary. Furthermore, the right half of Tab. V shows that the accuracy increases with model size. The 24M and 40M models correspond to using SVTR-B and SVTR-L [5] as the image encoder, with the remaining architecture keeping the same. We observe a similar large improvement gap (1.09% v.s. 0.09%) between the two methods, suggesting that attribute prediction instructions are the key to increase model performance. However, to fairly compare with other models, here we take SVTR-B as the backbone and IGTR models trained from 20 epochs for further analysis.
4.3 Zero-Shot Recognition
@@
As aforementioned, IGTR can evoke a deep understanding of the text image by training exclusively with attribute prediction instructions. It is observed that text recognition can also be achieved by using certain questions. We term such a pipeline as zero-shot recognition, i.e., IGTR-Zero. Specifically, we use two types of questions to enable text inference. They are constrainted frequency and search status. The quantitative results are given in Tab. VI and Fig. 5 illustrates the recognition procedures. The former question, by combining with another already learned edge char question that tells the first character, enables stepwise inference of the second (by setting the second variable in the question as 2), third, and subsequent characters, thus recognizing the whole text. The recognition process shows certain chain-of-thought capabilities as current large models [21, 72]. However, this inference chain is relatively complex and the model only achieves moderate success on challenging datasets (80.22% on Union14M-L). The latter one, by traversing all characters with all positions, presents a more straightforward and less chain-dependent way to determine the text. It reaches an accuracy of 82.33% on Union14M-L. The results indicate that the text image is well understood by instruction-based attribute learning.
@@ 4.4 Rare and Similar Character Recognition
Recognizing rarely appearing and morphologically similar characters are two typical challenges in STR, especially for Chinese. Owning to the flexibility of instruction-guided learning, we devise an elegant scheme to alleviate the two challenges by adjusting the rule of instruction sampling. The first part is for rare characters. Specifically, we first inspect the CTR dataset [59], characters with less than 50 occurrences in the validation set are treated as rare characters. When a text instance contains a rare character $c_{r}$, attributes related to $c_{r}$ are used to generate question-answer pairs, while other attributes are randomly divided into condition and question-answer pairs as before. In addition, for rare characters that are not in the input text, five of them are randomly selected and are used to generate questions of search status. For example, Is $c_{i}$ the j-th character in the image?, where $c_{i}$ is one randomly selected rare character. By adjusting the sampling rules as above, rare character attributes receive more attention, and thus learned better.
Then we explain how the rule is adjusted for morphologically similar characters. For a given character, we randomly select three characters, which have been misclassified as other characters, as its morphologically similar characters, or all characters are selected if the misclassified characters are less than three. Then, given a text instance, search status questions are constructed. For example, for a similar character pair “大-太”, where the first character is the appeared character and the second is its similar character. The answer would be Yes when the question is Is the $i$-th character “大”?, and No when the question is Is the $i$-th character “太”?. Morphologically similar characters will be better distinguished with the question-answer pairs above.
We term the adjustments above as targeted strengthening (TS) and devise a model termed IGTR-PR-TS to incorporate them. Results on the CTR dataset [59] are presented in Tab. VII and Tab. VIII. In Tab. VII, rare characters are grouped into three classes according to their occurrences. Compared to the raw IGTR-PR, the accuracy gaps are 3.59%, 1.68%, and 1.07% for characters appearing 1 to 10, 11 to 30, and 31 to 50 times, respectively. Larger gains are observed for severely rare characters. Meanwhile in Tab. VIII, each character pair gives two values, e.g., 8/5 for the “莱-菜” pair. It means that IGTR-PR classifies “莱” as “菜” 8 times, and the times is reduced from 8 to 5 when IGTR-PR-TS is employed. As can be seen, recognition errors occurred between morphologically similar characters are largely reduced. Both experiments convincingly verify the flexibility of IGTR and its potential to address typical STR challenges.
@@ 4.5 Comparison with State-of-the-Art
TABLE XI: Results on CTR dataset tested against existing models.
@@
We then compare IGTR with previous STR models on both Common and Union14M-L [58] benchmarks. The results trained based on synthetic and Union14M-L datasets are presented in Tab. IX and Tab. X, respectively. Meanwhile, we also train IGTR in another manner: first pre-training with attribute prediction instructions on synthetic datasets [49, 50, 51], and then fine-tuning with recognition instructions on Union14M-L. The obtained models are marked with a suffix PT, e.g., IGTR-PR-PT and IGTR-AR-PT.
@@
We first inspect the results trained based on synthetic datasets. As shown in Tab. IX, in Ours we give results of four IGTR models corresponding to the four recognition pipelines. Note that IGTR-PR-RI means RI is further incorporated into the PR pipeline. IGTR-AR ranks the top among 11 of the 13 evaluated subsets from both Common and Union14M-L benchmarks. It surpasses IGTR-PR by 0.48% on common benchmarks and 2.67% Union14M-L, respectively, demonstrating the effectiveness of encoding already recognized characters via condition, especially on more challenging Union14M-L. For IGTR-PR-RI v.s. IGTR-PR, 1% improvement is observed on Union14M-L, showing that the re-identification can correct misrecognition to some extent. Moreover, all four IGTR models perform better than all existing models in Tab. IX. For example, IGTR-AR outperforms LPV-B [7], one of the most competitive previous models, by 1.0% and 5.67% on Common and Union14M-L benchmarks, respectively. Those results convincingly indicate the superiority of the comprehension-first and recognition-next paradigm.
@@
We then look into the results on real-world benchmarks in Tab. X. It is seen that much higher accuracy is observed when the models are trained on Union14M-L, highlighting the importance of more realistic training data. Similarly, IGTR models excel all existing models in Tab. X. IGTR-PR-PT and IGTR-AR-PT, the two models mixed synthetic data-based pre-training and real-world data-based fine-tuning, perform considerably well. IGTR-AR-PT ranks the top among 11 of the 13 evaluated subsets. This can be explained as more training data is involved, and meanwhile, it follows an easy-to-hard learning procedure, i.e., first pre-trains on less challenging synthetic data and then fine-tunes on more difficult real-world data. For IGTR-AR-PT, the accuracy gap to SVTR-B, one of the best previous models in Tab. X is prominent 2.62% and 6.74%, respectively. The results again demonstrate the superiority of our instruction-guided learning. When looking back to the zero-shot recognition results, we observe that the results obtained based on the search status question (Tab. VI) are approaching even exceed existing top-performed models, especially on Union14M-L. This implies the effectiveness of the devised character attribute prediction instructions, which already endow the model with comprehensive text understanding. Note that all IGTR models only differ in training details. These models are all with a size of 24.1M, which is highly competitive compared to existing models.
@@
In Tab. XI, we also give the comparisons on Chinese benchmark [59]. IGTR still outperforms existing models by clear margins. When targeted strengthening (TS) is further employed we observe accuracy gains, which come from the better recognition of rarely appearing and morphologically similar characters. In addition, we observe that typical data augmentation is particularly useful for the CTR benchmark. Adding it can further improve the accuracy up to 3.80% for PR and 4.57% for AR. We argue that on the one hand, the image modality is relatively scarce when trained on the instruction-guided paradigm. On the other hand, the CTR dataset is not big enough for Chinese with thousands of character categories. Data augmentation can mitigate the two issues especially for difficult text, resulting in more prominent improvements on challenging subsets such as Scene and Handwriting. To sum, the experiments also validate the superiority of IGTR and its great cross-language generalization ability.

### v3

@@ 4.2 Ablations
Instruction variants. We conduct experiments to validate the effectiveness of the triplet-form instructions. The results in Tab. III include IGTR-PR (PR), IGTR-AR (AR) and IGTR-ER (ER), where the suffix PR means using the PR instruction in Tab. II for text inference and the others are defined similarly. In (a), only recognition instructions are utilized for IGTR training. In subsequent variants denoted by (b), (c), (d), and (j), we progressively add questions related to the referred attributes, where accuracy improvements are steadily obtained for all three models. The improvements reveal that the attributes all contribute positively, supporting that understanding diverse character attributes aids in recognition. Comparing the use of full instructions in (j) to (a), PR is improved by +1.94%, AR by +3.59%, and ER by +2.99%, indicating that AR and ER benefit more from the diverse questions. The results align with the fact that compared to PR, AR and ER own more recognition clues. They employ an iterative pipeline and can utilize previously recognized characters as linguistic clues. On the other hand, we perform another set of ablations on condition, starting from (e) with all questions but no condition ($cond$), and then progressively add a condition mode in (f)-(j), where the suffix is the condition index in Tab. I. IGTR enables AR and ER capabilities only when $cond$-3 and $cond$-4 are successively added, implying that enriching condition not only enhances accuracy but also enables different recognition pipelines. Additionally, by increasing K, the number of partitions, the accuracy steadily improved, again providing intuitive confirmation of the importance of improving instruction richness and diversity.
@@
CMFF components. As shown in Fig. 2, CMFF incorporates four cross-attention stages for inter-modal interaction. We devise variants by removing different stages, and the results are presented in Tab. IV. Removing any stage all results in a decrease in accuracy. For example, when query-to-condition and query-to-image stages are both removed, notable accuracy declines of 1.66% for PR, 3.63% for AR, and 2.95% for ER are observed. The results emphasize the necessity of incorporating both sufficient feature fusion stages and a single condition embedding, which serves as Key and Value for the query-to-condition stage. For AR and ER the condition is preceding decoded characters while for PR is set to None. This explains why such removal has a more pronounced impact on AR and ER, while PR is relatively less affected. Meanwhile, when question-to-condition and question-to-image stages are both removed, the accuracy declines are 1.58% for PR, 2.85% for AR, and 2.53% for ER, all preserving a consistent decreasing trend as above albeit with a smaller magnitude. This is because the two stages mainly focus on conducting finer feature interactions, and thus have relatively small affections on model performance. Furthermore, PR, AR, and ER report inference speeds of 3.98 ms, 10.3 ms and 9.52 ms, respectively. The speed differences align with their respective recognition pipelines. Note that they run faster than many previous models within their recognition pipelines. In addition, as shown in Tab. IV, removing the cross-attention stages only triggers minor speed acceleration. The results, in turn, verify that CMFF is quite efficient.
@@
TABLE VI: The recognition results of using different attribute prediction instructions.
@@
IGTR-PR 63.88 84.93 88.79 72.25
IGTR-PR-TS 67.47 86.61 89.86 75.05
TABLE VIII: Morphologically similar characters and mis-recognitions on CTR [59] using IGTR-PR and IGTR-PR-TS. Take “{CJK}UTF8gbsn莱-{CJK}UTF8gbsn菜” and 8/5 as the example. It means that without considering TS, “{CJK}UTF8gbsn莱” has been recognized as “{CJK}UTF8gbsn菜” for 8 times, and the number is reduced to 5 when TS is employed.
@@
TABLE X: Results on English benchmarks tested against existing models when trained on real-world Union14M-L training set. denotes that the result is obtained by training the model on Union14M-L using the code they released.
@@
Scalability of IGTR. We conduct ablations on training data volume and model size. Specifically, we select IGTR-PR and vanilla-PR, where the latter is trained solely using the PR instruction in Tab. II. The results are presented in the left half of Tab. V, where different training epochs are considered. We observe a notable accuracy gain of 2.78% when increasing epochs from 20 to 60 for IGTR-PR. This improvement is accompanied by the expansion of training data volume, as IGTR-PR samples attribute prediction instructions individually at each epoch. The trained model benefits from richer instructions. In contrast, the improvement of vanilla-PR is 0.91%, solely from more training epochs that generate a slightly better model optimization. The significant improvement gap (2.78% v.s. 0.91%) implies that our instruction-guided training is more flexible and can produce more powerful STR models when necessary. Furthermore, the right half of Tab. V shows that the accuracy increases with model size. The 24M and 40M models correspond to using SVTR-B and SVTR-L [5] as the image encoder, with the remaining architecture keeping the same. We observe a similar large improvement gap (1.09% v.s. 0.09%) between the two methods, suggesting that diverse attribute prediction instructions have a greater potential to enhance the capability of the visual backbone. However, to fairly compare with other models, we take SVTR-B as the backbone and train IGTR models with fixed 20 epochs in subsequent evaluations.
4.3 Recognition by Attribute Prediction Instructions
@@
It is observed that the attribute prediction instructions, when intelligently combined, can also infer the whole text. We refer to this pipeline as attribute prediction-based recognition. To validate this pipeline, we first train the model on attribute prediction instructions, then employ two questions to enable text inference. They are constrainted frequency and search status. The quantitative results are given in Tab. VI, and Fig. 5 illustrates the recognition procedures. The former question, when combined with another already learned edge char question that tells the first character, enables a step-by-step inference of subsequent characters, thus recognizing all characters. Specifically, we first set the second variable in question constrainted frequency to 2 and enumerate all the characters, from which the second character can be determined. Then, the third character can be deduced as similar by modifying the second variable in the question to 3, and so on. However, this inference chain is relatively complex and the model only achieves moderate success on challenging datasets (80.22% on Union14M-L). The latter question, by traversing all characters across all positions, presents a more straightforward and less chain-dependent way to determine the text. It reaches an accuracy of 82.33% on Union14M-L. The results demonstrate that by training solely on attribute prediction instructions, without the inclusion of text recognition instructions, IGTR can inherently develop certain character-level recognition capabilities. The text is then recognized when questions in attribute prediction instructions are appropriately utilized.
@@ 4.4 Rare and Similar Character Recognition
Recognizing rarely appearing and morphologically similar characters are two typical challenges in STR, especially for Chinese which has a large character vocabulary. Owning to the flexibility of instruction-guided learning, we devise an elegant scheme to alleviate the two challenges by adding certain rules to instruction sampling. The first part is for rare characters. Specifically, we first inspect the CTR dataset [59], characters with less than 50 occurrences in the validation set are treated as rare characters. When a text instance contains a rare character $c_{r}$, attributes related to $c_{r}$ are used to generate question-answer pairs, while other attributes are randomly divided into condition and question-answer pairs as before. In addition, for rare characters that are not in the input text, five of them are randomly selected and are used to generate questions of search status. For example, Is $c_{i}$ the j-th character in the image?, where $c_{i}$ is one selected rare character. By adjusting the sampling rules as above, attributes related to rare characters receive more attention, and thus learned better.
Then we explain how the rule is adjusted for morphologically similar characters. For a given character, we randomly select three characters, which the given character has the most frequently been misclassified to, as its morphologically similar characters, or all characters are selected if the misclassified characters are less than three. Then, given a text instance, specific search status questions are sampled for appeared characters in the form of these similar character pairs. For example, for a similar character pair “大-太”, where the first character is the appeared character and the second is its morphologically similar character. The answer is Yes when the question is Is the $i$-th character “大”?, and No when the question is Is the $i$-th character “太”?. By reinforcing these facts during model training, morphologically similar characters can be better distinguished.
We term the two adjustments above as Targeted Strengthening (TS) and devise a model termed IGTR-PR-TS to incorporate them. Results on the CTR dataset [59] are presented in Tab. VII and Tab. VIII, respectively. In Tab. VII, rare characters are grouped into three classes according to their occurrences. Compared to the raw IGTR-PR, the accuracy improvements are 3.59%, 1.68%, and 1.07% for characters appearing 1 to 10, 11 to 30, and 31 to 50 times, respectively. Larger gains are observed for severely rare characters. Meanwhile in Tab. VIII, each character pair gives two values, e.g., 8/5 for the “莱-菜” pair. It means that IGTR-PR classifies “莱” as “菜” 8 times, and the times is reduced from 8 to 5 when IGTR-PR-TS is employed. As can be seen, recognition errors occurred between morphologically similar characters are largely reduced. Both experiments convincingly verify the flexibility of IGTR and its potential to address typical STR challenges.
@@ 4.5 Comparison with State-of-the-Art
TABLE XI: Results on the Occluded Scene Text dataset [29]. All the models are trained on Union14M-L training set except for those suffixed with PT.
Vanilla-PR IGTR-PR IGTR-AR IGTR-PR-PT
WOST 82.5 83.2 85.1 89.7
HOST 64.1 65.3 67.6 82.1
Avg 73.30 74.25 76.32 85.90
IGTR-AR-PT SVTR-B [5] LPV-B [7] MAERec [52]
WOST 90.2 80.6 80.7 79.1
HOST 82.7 65.0 65.9 65.2
Avg 86.45 72.8 73.30 72.15
TABLE XII: Evaluations on pre-training methods, where the attribute prediction-based pre-training is substituted by three existing STR pre-training schemes. Com and U14M denote Common and Union14M-L benchmarks, respectively.
Pretraining Methods SeqCLR [78] MAERec [52] DiG [70]
Com U14M Com U14M Com U14M
IGTR-PR 96.36 83.73 96.60 84.34 96.51 84.91
IGTR-AR 96.71 85.36 96.85 86.30 96.90 86.74
TABLE XIII: Results on CTR dataset tested against existing models.
@@
We compare IGTR with previous STR models on both Common and Union14M-L benchmarks. The results trained based on synthetic and Union14M-L training set are presented in Tab. IX and Tab. X, respectively. Meanwhile, we also train IGTR in another manner: first pre-trained with attribute prediction instructions on the synthetic datasets, and then fine-tuned with text recognition instructions on Union14M-L training set. The obtained models are marked with a suffix PT, e.g., IGTR-PR-PT and IGTR-AR-PT.
@@
We first inspect the results trained on synthetic datasets. As shown in Tab. IX, in Ours we give results of four IGTR models corresponding to the four text recognition pipelines. Note that IGTR-PR-RI means RI is further incorporated into the IGTR-PR pipeline. IGTR-AR ranks the top among 11 of the 13 evaluated subsets from both Common and Union14M-L benchmarks. It surpasses IGTR-PR by 0.48% on Common and 2.67% on Union14M-L in terms of accuracy, respectively, demonstrating the effectiveness of seeing already decoded characters during recognition, especially on more challenging Union14M-L. For IGTR-PR-RI v.s. IGTR-PR, 1% improvement is observed on Union14M-L, showing that the re-identification can correct misrecognition to some extent. Moreover, the four IGTR models perform almost better than all the compared existing models in Tab. IX. For example, IGTR-AR outperforms LPV-B [7], one of the most competitive previous models, by 1.0% and 5.67% on Common and Union14M-L, respectively. Those results convincingly indicate the superiority of our comprehension-first and recognition-next paradigm.
@@
We then examine the models trained on real-world datasets. The results on Union14M-L are depicted in Tab. X. All models show a significant improvement in accuracy when trained on Union14M-L training set, underscoring the importance of real-world training data. Similarly, IGTR models excel all existing models in Tab. X except for OTE [75] in a few subsets within Common. IGTR-PR-PT and IGTR-AR-PT, the two models mixed synthetic data-based pre-training and real-world data-based fine-tuning, perform considerably well. IGTR-AR-PT achieves the highest rank in 11 out of the 13 evaluated subsets. The results showcase that the first pre-training and then fine-tuning scheme is a superior training paradigm. This can be explained as more training data is involved, and meanwhile, this paradigm follows an easy-to-hard learning procedure, i.e., first pre-trains on less challenging synthetic data and then fine-tunes on more difficult real-world data. For IGTR-AR-PT, the accuracy gaps to SVTR-B, one of the best previous models in Tab. X, are prominent 2.62% on Common and 6.74% on Union14M-L, respectively. The results again demonstrate the superiority of our instruction-guided learning. When looking back to the attribute prediction-based recognition results in Tab. VI, we observe that the results obtained based on the search status question already surpass existing top-performed models on Union14M-L. This implies the effectiveness of the devised character attribute prediction instructions, which already endow the model with comprehensive text understanding. Note that all IGTR models only differ in training details. These models are all with a model size of 24.1M, which is highly competitive compared to existing models.
We further analyze the performance of IGTR in Contextless, a subset in Union14M-L containing text that has no semantic meaning and is not in the dictionary. The results are listed in Tab. X and from which two observations can be inferred. First, IGTR models still get quite impressive accuracy. The accuracy of IGTR-PR-PT reaches 83.6% while IGTR-AR-PT achieves 83.4%. Both results surpass previous leading methods such as LPV-B [7] and MAERec [52]. Since contextless text provides limited linguistic prior, the accuracy gains are mainly because the rich and diverse instructions compel the model to learn character attributes from different aspects during IGTR training. Thus fine-grained visual content has been well understood and robust visual recognition capability is established. Second, IGTR-PR-PT performs slightly better than IGTR-AR-PT, which is opposed to the comparison on the other six subsets in Union14M-L. This implies that the availability of previously decoded characters in the AR pipeline does not contribute to the recognition, which also aligns with the text’s contextless nature. Both observations indicate IGTR’s superior text recognition capability, and IGTR-PR-PT and IGTR-RR-PT comprehensively understand character attributes after the pre-training.
To further assess the robustness of IGTR models, we conduct experiments on Occluded Scene Text (OST) dataset [29], which includes the Weakly Occluded Scene Text (WOST) dataset and the Heavily Occluded Scene Text (HOST) dataset. The two datasets pose a unique challenge in that a portion of text is obscured, therefore accurately recognizing the text requires the model to have certain text reasoning capability. As shown in Tab. XI, all IGTR models exhibit superiority over the compared methods, i.e., SVTR-B [5], LPV-B [7] and MAERec [52], indicating the advantages of instruction-guided learning in augmenting the capture of linguistic association. Interestingly, IGTR-PR-PT and IGTR-AR-PT achieve much higher accuracy compared to the IGTR-based competitors, especially in the more challenging HOST dataset. This observation suggests that the amalgamation of increased training data with instructional guidance effectively models the linguistic context of the text, and leading to enhanced performance.
@@
In Tab. XIII, we also give the results on CTR [59], a Chinese benchmark. IGTR models still outperform existing methods by clear margins. Further accuracy gains are observed when targeted strengthening (TS) is employed. The improvements come from the better recognition of rarely appearing and morphologically similar characters. In addition, we observe that data augmentation is particularly useful for the CTR benchmark. Adding the augmentation can further improve the accuracy up to 3.80% for IGTR-PR-TS and 4.57% for IGTR-AR-TS. We argue that, on one hand, the image modality constitutes a minority in terms of quantity when trained following the instruction-guided paradigm. On the other hand, the CTR dataset is not big enough for Chinese with thousands of character categories. Data augmentation can mitigate the two issues especially for difficult text, resulting in more prominent improvements on challenging subsets such as Scene and Handwriting. To sum, the experiments also validate the superiority of IGTR and its great cross-language generalization ability. Note that for Chinese the model size increases moderately due to the incorporation of a larger character vocabulary compared to English.
4.6 Distinction with existing pre-training methods
Previous methods introduced several general pre-training techniques into STR. They can be roughly categorized into three types: contrastive learning-based methods like SeqCLR [78] and PerSec [65], mask auto-encoder-based methods like MAERec [52], and methods like DiG [70] that combines both techniques. We select the open-sourced SeqCLR, MAERec, and DiG from the three types and compare our IGTR-PR-PT and IGTR-AR-PT with them. For the three existing methods, IGTR architecture is first pre-trained on Union14M-U [52], an unlabeled real-world dataset containing approximately 10 million samples. Afterwards, the results in Tab. XII are obtained by fine-tuning on Union14M-L training set. Since IGTR-PR-PT and IGTR-AR-PT are both pre-trained on automatically rendered synthetic datasets, all the compared methods do not require manual annotation and their fine-tuning step is the same. The two training protocols constitute a fair comparison among the methods.
As depicted in Tab. XII, models pre-trained based on DiG get slightly better results than those pre-trained based on SeqCLR and MAERec, highlighting the merit of combining both contrastive learning-based and mask auto-encoder-based pre-training. Moreover, all three pre-training methods positively contribute to training better recognition models, where performance gains are consistently observed when comparing the raw IGTR-PR and IGTR-AR results in X. The results confirm the effectiveness of using pre-training techniques. Moreover, IGTR-PR-PT outperforms DiG-based IGTR-PR by 0.69% in Common and by 2.01% in Union14M-L. Similarly, IGTR-AR-PT surpasses DiG-based IGTR-AR by 0.44% in Common and by 0.91% in Union14M-L. The results indicate that our attribute prediction-based pre-training provides more beneficial knowledge for text recognition tasks compared to the three pre-training methods, even when only trained on synthetic datasets. We posit that this superiority arises from the diversity of our instructions, which leads to a more comprehensive pre-training regime and consequently better model training. This reaffirms the significance of understanding diverse character attributes.